40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 21     Entry time: Mon Jan 28 18:02:04 2008
Author: Stefan Ballmer 
Type: Laser 
Category: General 
Subject: 1.2MHz Laser intensity oscillation 
While working on the PDH loop for the PMC I noticed
a 1.2MHz oscillation on the laser intensity:

- It seems to be a bistable state: it was generally
started by some transient in the piezo input of the laser.
- The oscillation persists even when all control inputs to the laser
(EOM, PZT, AOM) are disconnected.
- It's extremely strong: about 20% fractional modulation
- Power-cycling the laser usually fixes the problem.
ELOG V3.1.3-